OpenStudy (anonymous):

Solve log x2 + 1 = 5. Round to the nearest thousandth if necessary.

OpenStudy (anonymous):

log(x^2) = 2logx

OpenStudy (anonymous):

What would x equal?

OpenStudy (anonymous):

using the fact that \[\log_ab=x\] can also be written as \[a^x=b\] \[\log x^2 +1 =5\]

OpenStudy (anonymous):

is it \[\log (x^2+1)=5\]or \[\log (x^2) + 1 =5\]??

OpenStudy (anonymous):

The second one

OpenStudy (anonymous):

\[\log (x^2)+1=5\] \[\log (x^2)=4\] using the fact i gave above..^ ^ \[x^2=10^4\] \[x=\sqrt{10^4}\] \[x=100\]

OpenStudy (anonymous):

Thank you so much for showing me how to solve the problem!